OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (8, 1) (7, 9) i got y=-64+1x?

OpenStudy (anonymous):

aroub can you answer my question please

OpenStudy (anonymous):

i need your help on this one

OpenStudy (aroub):

(8,1) (7,9) =-8 y=mx+b (8,1) 1=-8(8)+b , 1=-64+b , b=65 y=-8x+65

OpenStudy (mathteacher1729):

You should quickly be able to tell that the slope is negative because the difference in the x coordinates is only one, but the difference in the y coordinates is -8. (see attached).

OpenStudy (anonymous):

thanks!